An oil company bores a hole 80ksh deep estimate the cost of boring the hole if the cost is 30ksh for drilling the first meter with an increase in cost of 2ksh per meter for each succeeding meters
Found 2 solutions by Boreal, greenestamps:
Answer by Boreal(15235)

this is a sum of an arithmetic series and I assume the hole is 80 m deep, but ksh is a unit of cost and not of depth. It that correct? Or does someone want to know how deep they can drill for 80 ksh? I will do the first one.
a1=30
n=80
d=2
sum=(80/2)(2a1+(n-1)d))
=40(60+79*2)
=40(218)
=8720 ksh

Answer by greenestamps(13200)



Cost of 1st meter: 30
Cost of 80th meter: 30+79(2)=30+158=188

Average cost per meter (= average cost of first and last meter, because the sequence is arithmetic): (30+188)/2 = 218/2 = 109

The total cost is the average cost per meter, multiplied by the number of meters.

109*80=8720

ANSWER: 8720ksh